# More is better: Strong constraints on the stellar properties of LEGA-C galaxies with Prospector

### v1

@@ Abstract
Methods. We estimate the galaxy properties with the Bayesian spectral energy distribution (SED) framework Prospector. We fit spectroscopy and broadband photometry together, drawn from the LEGA-C DR3 and UltraVISTA catalogs respectively.
Results. We report a positive correlation between light-weighted ages and stellar velocity dispersion (). The trend with is weaker for the mass-weighted ages and stellar metallicity (). At fixed , we find a tentative correlation between and stellar age. On average, quiescent galaxies are characterized by high , they are Gyr older, less dusty, with steeper dust attenuation slopes (due to a lower optical depth) compared to star-forming galaxies. Conversely, star-forming galaxies are characterized by significantly higher dust optical depths and shallower (grayer) attenuation slopes. Low mass (high mass) star-forming galaxies have lower (higher) , while their stellar populations are on average younger (older). A key pragmatic result of our study is that a linear-space metallicity prior is preferable to a logarithmic-space one when using photometry alone, as the latter biases the posteriors downward.
Conclusions. Spectroscopy greatly improves stellar population measurements and is required to provide meaningful constraints on age, metallicity, and other properties. Pairing spectroscopy with photometry helps resolving the dust–age–metallicity degeneracy. Spectroscopic data yield more accurate mass- and light-weighted ages, with ages inferred from photometry alone suffering such large uncertainties that their utility is limited. Stellar metallicities are constrained by our spectroscopy, but precise measurements remain challenging (and impossible with photometry alone), particularly in the absence of Mg and Fe lines redward of 5000 Å in the observed spectrum.
@@
Determining the properties of the stellar populations in galaxies is essential to describe how galaxies assemble their mass over cosmic time. The simplest method to infer these properties and whether the process of star formation is ongoing or not, is through the analysis of broadband photometry. A common practice is to combine the light of different optical and near-infrared (NIR) bands (galaxy colors) to determine the stellar population properties (e.g. Strateva et al. 2001; Baldry et al. 2004; Bell et al. 2004; Labbé et al. 2005; Faber et al. 2007; Wuyts et al. 2007; Williams et al. 2009; Salim 2014). While optical colors and multiwavelength photometry (both broadband and narrowband) are useful for studying galaxies, they are insufficient on their own to fully constrain a galaxy’s chemical composition and star-formation history (SFH) (e.g. Chaves-Montero & Hearin 2020; Nersesian et al. 2024; Csizi et al. 2024).
@@
It is well known that inferences from galaxy photometry suffer strongly from the age-dust-metallicity degeneracies (e.g. Worthey et al. 1994; Silva et al. 1998; Devriendt et al. 1999; Pozzetti & Mannucci 2000; Bell & de Jong 2001; Walcher et al. 2011). One way to mitigate these degeneracies, and rigorously determine the SFHs and chemical composition of galaxies, is by complementing the photometric SEDs with spectroscopic information (e.g. Worthey et al. 1994; Bruzual & Charlot 2003; Trager et al. 2000a; Gallazzi & Bell 2009). In fact, the importance of optical spectroscopy on the study of stellar populations was established early on, with the measurements of the strengths of specific absorption-line features in nearby galaxies (e.g. using the Lick index system, Worthey et al. 1994). The use of the Lick index system allowed the constraints on the mean stellar ages and metallicities of galaxies (e.g. Jørgensen 1999; Trager et al. 2000b, a; Kuntschner 2000; Kauffmann et al. 2003a; Gallazzi et al. 2006; McDermid et al. 2015; Wu et al. 2018b). The recent studies by Johnson et al. (2021) and Tacchella et al. (2022) also highlighted the importance of spectroscopy to constrain the stellar parameters, along with the necessity of both spectroscopy and photometry to alleviate the dust–age–metallicity degeneracies.
@@
Advancements in the development of stellar population synthesis (SPS) models based on high spectral resolution stellar spectra, both empirical and theoretical, allow the production of highly resolved synthetic galaxy spectra. Some examples of these SPS models that incorporate the latest developments in stellar evolution theory include: GALAXEV (Bruzual & Charlot 2003), Maraston (2005); Maraston & Strömbäck (2011); Maraston et al. (2020), the Flexible Stellar Populations Synthesis (FSPS) code (Conroy et al. 2009; Conroy & Gunn 2010), and BPASS (Eldridge et al. 2017; Stanway & Eldridge 2018; Byrne et al. 2022). By comparing and matching the observed spectra with these highly complex SPS models, it is possible to characterize the relevant stellar properties of galaxies including the metallicity and age (either mass-weighted or light-weighted) of the stellar populations. An accurate estimation of these two properties sets the stage for an accurate description of the chemical enrichment and SFH of galaxies.
The emergence of large photometric and spectroscopic surveys both at low-redshift (e.g. SDSS, York et al. 2000; GAMA, Driver et al. 2011), as well as at intermediate-redshift (e.g. COSMOS, Scoville et al. 2007; Weaver et al. 2022; 3D-HST, Brammer et al. 2012; Momcheva et al. 2016; MOSDEF, Kriek et al. 2015; VIPERS, Scodeggio et al. 2018), in combination with the high-spectral resolution SPS models (see review by Conroy 2013), led to the development of full-spectrum fitting tools (see reviews by Walcher et al. 2011; Pacifici et al. 2023).
@@
In the last decade, state-of-the-art spectral energy distribution (SED) tools have taken up the role of estimating the integrated properties of galaxies. A few examples among the most recently developed SED fitting frameworks are pPXF (Cappellari & Emsellem 2004; Cappellari 2017), STARLIGHT (Cid Fernandes et al. 2005), STECMAP (Ocvirk et al. 2006), VESPA (Tojeiro et al. 2007), MAGPHYS (da Cunha et al. 2008), FIT 3D (Sánchez et al. 2016; Lacerda et al. 2022), BEAGLE (Chevallard & Charlot 2016), FIREFLY (Wilkinson et al. 2017), FADO (Gomes & Papaderos 2017), Dense Basis (Iyer & Gawiser 2017; Iyer et al. 2019), BAGPIPES (Carnall et al. 2018), CIGALE (Boquien et al. 2019), PROSPECT, (Robotham et al. 2020), Prospector (Leja et al. 2017; Johnson et al. 2021), and piXedfit (Abdurro’uf et al. 2021). Different SED fitting codes offer different capabilities such as (1) the analysis of multiwavelength photometry from ground-based and space observatories, (2) the analysis of optical spectra either from slit or integral-field spectroscopic surveys, and (3) the combining of photometry and spectroscopy with a calibration model.
@@
In the local Universe, high-quality, and even spatially-resolved, optical spectra are easy to obtain (e.g. ATLAS3D, Cappellari et al. 2011; CALIFA, Sánchez et al. 2013; SAMI, Bryant et al. 2015; MaNGA, Bundy et al. 2015). However, obtaining high-quality spectroscopy for galaxies at higher redshifts is often more challenging, as such surveys must balance between achieving a high signal-to-noise ratio (S/N) and covering a large sample size. In addition, with increasing redshift the contribution of the sky background becomes more dominant, essentially dimming the surface brightness of galaxies (Hogg 1999). To account for this, many spectroscopic surveys targeted only bright emission lines originating from the ionized gas in galaxies.
@@
The advent of the Large Early Galaxy Astrophysics Census (LEGA-C; van der Wel et al. 2016, 2021) survey has changed the landscape of intermediate-redshift spectroscopy. The LEGA-C survey is an exceptional dataset that contains about 4,000 high S/N restframe optical spectra at observed redshift (or at a lookback time of Gyr). The galaxy sample was -band selected from UltraVista targets in the COSMOS field. So far, the analysis of the LEGA-C spectra has produced many interesting results, from the formation timescales and evolution pathways of galaxies at observed redshift (Chauke et al. 2018; Wu et al. 2018a; Kaushal et al. 2024), to the stellar kinematics (Bezanson et al. 2018; Straatman et al. 2018; Cole et al. 2020; de Graaff et al. 2021; D’Eugenio et al. 2023a, b; Barrientos Acevedo et al. 2023), and the interstellar medium (ISM) properties (Spilker et al. 2018; Barišić et al. 2020; Lewis et al. 2023).
@@
Spectroscopic analyses of the stellar properties of galaxies at a significant redshift have been limited to a small number of targets (e.g., Choi et al. 2014; Gallazzi et al. 2014; Belli et al. 2015). The LEGA-C survey, with its large statistical sample, offers a valuable opportunity to gain deeper insights into the stellar properties of galaxies during an epoch ( Gyr ago), when nearly 50% of all stars in the Universe had formed since (e.g. Ilbert et al. 2010; Muzzin et al. 2013a), and galaxies likely hosted different stellar populations.
@@
This paper is structured as follows. In Section 2 we describe the datasets we used and the properties of our galaxy sample. In Section 3 we present the physical model that was adopted within the Prospector inference framework to describe the spectrophotometric SEDs of the LEGA-C galaxies. In Section 4 we present the final products of SED fitting with Prospector, and discuss the underlying scaling relations between age, metallicity and dust in the the and SFR– diagrams. Finally, in Section 5 we summarize our key findings and conclusions.
@@ 2.1 Spectroscopy
The now retired VIMOS spectrograph (Le Fèvre et al. 2003) at the ESO Very Large Telescope (VLT) was used to obtain the spectroscopic observations of LEGA-C over the course of four years. The typical observed wavelength range of LEGA-C is 6300 Å8800 Å or restframe 3000 Å 5550 Å, with an effective spectral resolution of . The optical spectrum contains a significant number of absorption line features that show small galaxy-to-galaxy variations in terms of flux density. Many of these absorption features are often too weak to be accurately measured, and their detection may be further compromised by various systematic effects. In particular, a bias can be introduced in the the equivalent width measurement of the absorption line indices, if the variance of the spectrum is considered or not. Other systematic effects can be introduced during the steps of sky subtraction, noise modeling, and wavelength calibration. In van der Wel et al. (2021), a comprehensive description is given of the steps taken to minimize systematic effects and biases, by employing an approximately bias-free methodology. Along with the reduced LEGA-C spectra, a catalog was released with the Lick indices of 20 spectral absorption features, corrected for emission. In Gallazzi et al. (in prep.), an updated catalog will be released, increasing the number of objects with Lick index measurements, and effectively reducing the uncertainties of the measurements by a factor of 1.3.
@@ 2.3 Primary working galaxy sample
In our analysis, we just use the LEGA-C galaxies with a PRIMARY flag from the DR3 catalog. The PRIMARY flag indicates whether a galaxy was selected from the -band selected parent sample (1) or as a filler (0) (van der Wel et al. 2021). Our final working sample contains 2908 galaxies in the redshift range . We will hereafter refer to this sample of 2908 galaxies as primary galaxy sample. Figure 1 shows the spectrophotometric SEDs of these 2908 targets sorted with increasing redshift.
@@ 3 SED fitting with Prospector
In this section, we present the physical model that was adopted to describe the spectrophotometric SEDs of the LEGA-C galaxies. We also perform several quality tests of the fits. The physical model was generated within the Prospector inference framework (Johnson et al. 2021). Prospector is an SED fitting code that adopts a Bayesian forward modeling and Monte Carlo sampling of the parameter space. Prospector can generate gridless ‘on-the-fly’ SEDs by combining stellar, nebular, and dust models into composite stellar populations. One of the strong capabilities of Prospector is the treatment of photometric and spectroscopic data using a flexible spectrophotometric calibration. Another advantage of Prospector is the availability of nonparametric SFHs with various prior distributions and parameterization schemes. The benefit of using a nonparametric SFH is that we impose a much weaker prior on the shape of the functional form of the SFH. It has also been shown that a nonparametric SFH can outperform the traditional parametric SFH by vastly improving the stellar mass estimates (Lower et al. 2020). A summary of the free parameters and range of the prior distributions of our physical model is given in Table 1. Our physical model largely follows the model presented in Leja et al. (2017, 2018, 2019b). Lastly, we used the nested sampler dynesty (Skilling 2004, 2006; Koposov et al. 2022) to simultaneously estimate both the Bayesian evidence and the posterior distributions. dynesty also allows the dynamic sampling of the parameter space to maximize a chosen objective function as the fit proceeds.
@@ 3.2.1 Reduced distribution
Here represents the flux densities of the observed spectra, are the model spectra, are the observed uncertainties, and is the degrees of freedom (dof), calculated as the wavelength elements (i.e. ) minus the free parameters in our model (i.e. 20 free parameters). The distribution is shown in Fig. 2. We find that the peak of the distribution and the median value of the histogram are at the same value . The percentile is at a value of 0.82 and the is at 1.77. Out of the 2908 modeled spectra, 88% have a . About 11% of the modeled spectra have a , while only a very small fraction of 1% have . Overall, we find an excellent agreement between the observed and the modeled spectra. However, the prominent tail of high- values may suggest that these fits are dominated by systematic uncertainties.

### v2

@@ Abstract
Methods. We estimated the galaxy properties with the Bayesian spectral energy distribution (SED) framework Prospector. We fit spectroscopy and broadband photometry together, drawn from the LEGA-C DR3 and UltraVISTA catalogs, respectively.
Results. We report a positive correlation between light-weighted ages and stellar velocity dispersion (). The trend with is weaker for the mass-weighted ages and stellar metallicity (). At fixed , we find a tentative correlation between and stellar age. On average, quiescent galaxies are characterized by high ; they are Gyr older, less dusty, and have steeper dust attenuation slopes (due to a lower optical depth) compared to star-forming galaxies. Conversely, star-forming galaxies are characterized by significantly higher dust optical depths and shallower (grayer) attenuation slopes. Low-mass (high-mass) star-forming galaxies have lower (higher) , while their stellar populations are on average younger (older). A key pragmatic result of our study is that a linear-space metallicity prior is preferable to a logarithmic-space one when using photometry alone, as the latter biases the posteriors downward.
Conclusions. Spectroscopy greatly improves stellar population measurements and is required to provide meaningful constraints on age, metallicity, and other properties. Pairing spectroscopy with photometry helps to resolve the dust–age–metallicity degeneracy. Spectroscopic data yield more accurate mass- and light-weighted ages, with ages inferred from photometry alone suffering such large uncertainties that their utility is limited. Stellar metallicities are constrained by our spectroscopy, but precise measurements remain challenging (and impossible with photometry alone), particularly in the absence of Mg and Fe lines redward of 5000 Å in the observed spectrum.
@@
Determining the properties of the stellar populations in galaxies is essential to describe how galaxies assemble their mass over cosmic time. The simplest method to infer these properties and determine whether the process of star formation is ongoing or not is through the analysis of broadband photometry. A common practice is to combine the light of different optical and near-infrared (NIR) bands (galaxy colors) to determine the stellar population properties (e.g., Strateva et al. 2001; Baldry et al. 2004; Bell et al. 2004; Labbé et al. 2005; Faber et al. 2007; Wuyts et al. 2007; Williams et al. 2009; Salim 2014). While optical colors and multiwavelength photometry (both broadband and narrowband) are useful for studying galaxies, they are insufficient on their own to fully constrain a galaxy’s chemical composition and star-formation history (SFH) (e.g., Chaves-Montero & Hearin 2020; Nersesian et al. 2024; Csizi et al. 2024).
@@
It is well known that inferences from galaxy photometry suffer strongly from the age-dust-metallicity degeneracies (e.g., Worthey et al. 1994; Silva et al. 1998; Devriendt et al. 1999; Pozzetti & Mannucci 2000; Bell & de Jong 2001; Walcher et al. 2011). One way to mitigate these degeneracies, and rigorously determine the SFHs and chemical composition of galaxies, is by complementing the photometric spectral energy distributions (SEDs) with spectroscopic information (e.g., Worthey et al. 1994; Bruzual & Charlot 2003; Trager et al. 2000a; Gallazzi & Bell 2009). In fact, the importance of optical spectroscopy on the study of stellar populations was established early on, with the measurements of the strengths of specific absorption-line features in nearby galaxies (e.g., using the Lick index system, Worthey et al. 1994). The use of the Lick index system allowed the constraints on the mean stellar ages and metallicities of galaxies (e.g., Jørgensen 1999; Trager et al. 2000b, a; Kuntschner 2000; Kauffmann et al. 2003a; Gallazzi et al. 2006; McDermid et al. 2015; Wu et al. 2018b). The recent studies by Johnson et al. (2021) and Tacchella et al. (2022) also highlight the importance of spectroscopy to constrain the stellar parameters, along with the necessity of both spectroscopy and photometry to alleviate the dust–age–metallicity degeneracies.
@@
Advancements in the development of stellar population synthesis (SPS) models based on high spectral resolution stellar spectra, both empirical and theoretical, allow for the production of highly resolved synthetic galaxy spectra. Some examples of these SPS models that incorporate the latest developments in stellar evolution theory include: GALAXEV (Bruzual & Charlot 2003), Maraston (2005); Maraston & Strömbäck (2011); Maraston et al. (2020), the Flexible Stellar Populations Synthesis (FSPS) code (Conroy et al. 2009; Conroy & Gunn 2010), and BPASS (Eldridge et al. 2017; Stanway & Eldridge 2018; Byrne et al. 2022). By comparing and matching the observed spectra with these highly complex SPS models, it is possible to characterize the relevant stellar properties of galaxies including the metallicity and age (either mass-weighted or light-weighted) of the stellar populations. An accurate estimation of these two properties sets the stage for an accurate description of the chemical enrichment and SFH of galaxies.
@@
The emergence of large photometric and spectroscopic surveys both at a low-redshift (e.g., SDSS, York et al. 2000; GAMA, Driver et al. 2011), as well as at intermediate-redshift (e.g., COSMOS, Scoville et al. 2007; Weaver et al. 2022; 3D-HST, Brammer et al. 2012; Momcheva et al. 2016; MOSDEF, Kriek et al. 2015; VIPERS, Scodeggio et al. 2018), in combination with the high-spectral resolution SPS models (see review by Conroy 2013), led to the development of full-spectrum fitting tools (see reviews by Walcher et al. 2011; Pacifici et al. 2023). These state-of-the-art SED fitting tools have taken up the role of estimating the integrated properties of galaxies. A few examples among the most recently developed SED fitting frameworks are pPXF (Cappellari & Emsellem 2004; Cappellari 2017), STARLIGHT (Cid Fernandes et al. 2005), STECMAP (Ocvirk et al. 2006), VESPA (Tojeiro et al. 2007), MAGPHYS (da Cunha et al. 2008), FIT 3D (Sánchez et al. 2016; Lacerda et al. 2022), BEAGLE (Chevallard & Charlot 2016), FIREFLY (Wilkinson et al. 2017), FADO (Gomes & Papaderos 2017), Dense Basis (Iyer & Gawiser 2017; Iyer et al. 2019), BAGPIPES (Carnall et al. 2018), CIGALE (Boquien et al. 2019), PROSPECT, (Robotham et al. 2020), Prospector (Leja et al. 2017; Johnson et al. 2021), and piXedfit (Abdurro’uf et al. 2021). Different SED fitting codes offer different capabilities such as being able to (1) analyze multiwavelength photometry from ground-based and space observatories, (2) analyze optical spectra either from slit or integral-field spectroscopic surveys, and (3) combine photometry and spectroscopy with a calibration model.
@@
In the local Universe, high-quality, and even spatially-resolved, optical spectra are easy to obtain (e.g., ATLAS3D, Cappellari et al. 2011; CALIFA, Sánchez et al. 2013; SAMI, Bryant et al. 2015; MaNGA, Bundy et al. 2015). However, obtaining high-quality spectroscopy for galaxies at higher redshifts is often more challenging, as such surveys must balance between achieving a high signal-to-noise ratio (S/N) and covering a large sample size. In addition, with increasing redshift the contribution of the sky background becomes more dominant, essentially dimming the surface brightness of galaxies (Hogg 1999). To account for this, many spectroscopic surveys targeted only bright emission lines originating from the ionized gas in galaxies.
@@
The advent of the Large Early Galaxy Astrophysics Census (LEGA-C; van der Wel et al. 2016, 2021) survey has changed the landscape of intermediate-redshift spectroscopy. The LEGA-C survey is an exceptional dataset that contains about 4,000 high S/N restframe optical spectra at observed redshift (or at a lookback time of Gyr). The galaxy sample was -band selected from UltraVista targets in the COSMOS field. So far, the analysis of the LEGA-C spectra has produced many interesting results, from the formation timescales and evolution pathways of galaxies at observed redshift (Chauke et al. 2018; Wu et al. 2018a; Kaushal et al. 2024), to the stellar kinematics (Bezanson et al. 2018; Straatman et al. 2018; Cole et al. 2020; de Graaff et al. 2021; D’Eugenio et al. 2023a, b; Barrientos Acevedo et al. 2023), and the interstellar medium (ISM) properties (Spilker et al. 2018; Barišić et al. 2020; Lewis et al. 2024).
@@
Spectroscopic analyses of the stellar properties of galaxies at a significant redshift have been limited to a small number of targets (e.g., Choi et al. 2014; Gallazzi et al. 2014; Belli et al. 2015). The LEGA-C survey, with its large statistical sample, offers a valuable opportunity to gain deeper insights into the stellar properties of galaxies during an epoch ( Gyr ago), when nearly 50% of all stars in the Universe had formed since (e.g., Ilbert et al. 2010; Muzzin et al. 2013a), and galaxies likely hosted different stellar populations.
@@
This paper is structured as follows. In Section 2 we describe the datasets we used and the properties of our galaxy sample. In Section 3 we present the physical model that was adopted within the Prospector inference framework to describe the spectrophotometric SEDs of the LEGA-C galaxies. In Section 4 we present the final products of SED fitting with Prospector, and discuss the underlying scaling relations between age, metallicity and dust in the and SFR– diagrams. Finally, in Section 5 we summarize our key findings and conclusions.
@@ 2.1 Spectroscopy
The now retired VIMOS spectrograph (Le Fèvre et al. 2003) at the ESO Very Large Telescope (VLT) was used to obtain the spectroscopic observations of LEGA-C over the course of four years. The typical observed wavelength range of LEGA-C is 6300 Å8800 Å or restframe 3000 Å 5550 Å, with an effective spectral resolution of . The optical spectrum contains a significant number of absorption line features that show small galaxy-to-galaxy variations in terms of flux density. Many of these absorption features are often too weak to be accurately measured, and their detection may be further compromised by various systematic effects. In particular, a bias can be introduced in the equivalent width measurement of the absorption line indices, if the variance of the spectrum is considered or not. Other systematic effects can be introduced during the steps of sky subtraction, noise modeling, and wavelength calibration. In van der Wel et al. (2021), a comprehensive description is given of the steps taken to minimize systematic effects and biases, by employing an approximately bias-free methodology. Along with the reduced LEGA-C spectra, a catalog was released with the Lick indices of 20 spectral absorption features, corrected for emission. In Gallazzi et al. (in prep.), an updated catalog will be released, increasing the number of objects with Lick index measurements, and effectively reducing the uncertainties of the measurements by a factor of 1.3.
@@ 2.3 Primary working galaxy sample
In our analysis, we just use the LEGA-C galaxies with a PRIMARY flag from the DR3 catalog. The PRIMARY flag indicates whether a galaxy was selected from the -band selected parent sample (1) or as a filler (0) (van der Wel et al. 2021). Our final working sample contains 2908 galaxies in the redshift range . Hereafter, we refer to this sample of 2908 galaxies as primary galaxy sample. Figure 1 shows the spectrophotometric SEDs of these 2908 targets sorted with increasing redshift.
@@ 3 SED fitting with Prospector
In this section, we present the physical model that was adopted to describe the spectrophotometric SEDs of the LEGA-C galaxies. We also perform several quality tests of the fits. The physical model was generated within the Prospector inference framework (Johnson et al. 2021). Prospector is an SED fitting code that adopts a Bayesian forward modeling and Monte Carlo sampling of the parameter space. Prospector can generate gridless ”on-the-fly” SEDs by combining stellar, nebular, and dust models into composite stellar populations. One of the strong capabilities of Prospector is the treatment of photometric and spectroscopic data using a flexible spectrophotometric calibration. Another advantage of Prospector is the availability of nonparametric SFHs with various prior distributions and parameterization schemes. The benefit of using a nonparametric SFH is that we impose a much weaker prior on the shape of the functional form of the SFH. It has also been shown that a nonparametric SFH can outperform the traditional parametric SFH by vastly improving the stellar mass estimates (Lower et al. 2020). A summary of the free parameters and range of the prior distributions of our physical model is given in Table 1. Our physical model largely follows the model presented in Leja et al. (2017, 2018, 2019b). Lastly, we used the nested sampler dynesty (Skilling 2004, 2006; Koposov et al. 2022) to simultaneously estimate both the Bayesian evidence and the posterior distributions. dynesty also allows the dynamic sampling of the parameter space to maximize a chosen objective function as the fit proceeds.
@@ 3.2.1 Reduced distribution
Here represents the flux densities of the observed spectra, are the model spectra, are the observed uncertainties, and is the degrees of freedom (dof), calculated as the wavelength elements (i.e., ) minus the free parameters in our model (i.e., 20 free parameters). The distribution is shown in Fig. 2. We find that the peak of the distribution and the median value of the histogram are at the same value . The percentile is at a value of 0.82 and the is at 1.77. Out of the 2908 modeled spectra, 88% have a . About 11% of the modeled spectra have a , while only a very small fraction of 1% have . Overall, we find an excellent agreement between the observed and the modeled spectra. However, the prominent tail of high- values may suggest that these fits are dominated by systematic uncertainties.